Suppose you save $4,000 per year at the end of each year for 30 years and earn 5% interest per year. How much will you have at the end of 30 years?


1
Expert's answer
2022-01-12T14:40:59-0500

Annual savings (P) = $4,000 per year

Annual interest rate (r) = 5.00% per year

Number of periods (n) = 30 years

Therefore, Future Value of an Ordinary Annuity "= P \times \frac{(1+ r)^n - 1}{ r }"

"= \$4,000 \times \frac{(1 + 0.05)^{30} - 1}{ 0.05} \\

= \$4,000 \times \frac{4.3219424 - 1}{ 0.05} \\

= \$4,000 \times \frac{3.3219424 }{0.05} \\

= \$4,000 \times 66.4388475 \\

= \$265,755.39"